OpenStudy (crashonce):

AP GP help @ganeshie8

OpenStudy (crashonce):

If the quadratic equation \[ax ^{2}-2bx+c=0\] has two equal roots, which of the following are true?

OpenStudy (crashonce):

1: both roots are b/a 2: a,b,c are GP 3. a,b,c are AP

ganeshie8 (ganeshie8):

Discriminant has to be 0 for a quadratic to have equal roots

ganeshie8 (ganeshie8):

find the discriminant of given equation and set it equal to 0

OpenStudy (crashonce):

2b^2-4ac = 0????

ganeshie8 (ganeshie8):

(2b)^2 - 4ac = 0

ganeshie8 (ganeshie8):

cancel 4, you get b^2 - ac = 0 b^2 = ac b = sqrt(ac) that means b is the geometric mean of a and c

ganeshie8 (ganeshie8):

that also means a, b, c are in GP

OpenStudy (crashonce):

um what about the roots

ganeshie8 (ganeshie8):

yes both roots will be b/a

OpenStudy (crashonce):

thaks

ganeshie8 (ganeshie8):

b/c when D = 0, quadratic formula gives : \[x = \dfrac{2b \pm 0}{2a} \]